Tom had left home and had ridden his bike for 7 km before his chain broke. Tom had to walk home after this. During his trip back, Tom estimated that he biked 5 times faster than he walked. If the entire trip took him 5 hours and 24 minutes, what were his walking and riding speeds?
Answer by ptaylor(1894)

Distance(d) equals rate(r) times time(t) or d=rt; t=d/r and r=d/t
Let r=Tom's rate of speed walking
Then 5r=his rate of speed biking (assuming Tom's estimate is correct)
distance walking = distance riding
Now we are told that the time walking plus time riding equals 5hr 24min.
Time walking=7/r
Time riding=7/5r
So our equation to solve is:
7/r+7/5r=5hr 24min or 27/5 hr multiply each term by 5r
35+7=27r combine like terms
42=27r divide both sides by 27
r=1.5555555556 mph------------------Tom's rate of speed walking
5r=5*1.56=7.777777778 mph--------------Tom's rate of speed biking

CK
7/(1.5555555556)+7/(7.777777778)=5.4
4.49999999+0.8999999=5.4
5.39999999~5.4

Hope this helps----ptaylor